OpenStudy (anonymous):

Evaluate 31 + 2x when x = 4. A. 33 B. 37 C. 39 D. 132

OpenStudy (thread252):

39

OpenStudy (anonymous):

Exit Evaluate 0.53n + 5 when n = 3 and when n = −3. A. 6.59 and 3.41 B. −6.59 and 3.41 C. −6.59 and −3.41 D. 6.59 and −3.41

OpenStudy (anonymous):

Evaluate 4a + 5b when a = 2 and b = −3.1. A. 6.5 B. 16.1 C. 21.5 D. −7.5

OpenStudy (anonymous):

Which expression indicates 10 subtracted from 9n? A. 10 − 9n B. 9n − 10 C. 10 + (−9n) D. 9 − 10n

OpenStudy (anonymous):

Diane purchase 12 handkerchiefs for a certain amount of money. Using m for the cost of one handkerchief, which expression illustrates the amount of money spent on all the handkerchiefs? A. 12 − m B. 12 + m C. 12m D.

OpenStudy (anonymous):

Car rental at Q.T. Rental is $22 per day plus an initial deposit of $36. Which expression shows how much it will cost to rent a car for d number of days? A. 22d − 36 B. 22 − 36d C. 36 − 22d D. 36 + 22d

OpenStudy (anonymous):

Exit Tammy purchases 16 pairs of socks for a certain amount of money. Using s for the cost of one pair of socks, which expression represents the total cost of all 16 pairs of socks? A. 16 − s B. 16 + s C. 16s D.

OpenStudy (thread252):

too many question, I QUIT

OpenStudy (anonymous):

Evaluate ab + 3c when a = –6, b = –2, and c = 5. A. 0 B. 3 C. 4 D. 27

OpenStudy (anonymous):

A) 6.59 and 3.41 This solution is possible without actually multiplying .53 by 3 or -3; instead, use approximation and estimation: one-half of 3 is one and one-half; one -half of negative three is negative 1.5

OpenStudy (anonymous):

Evaluate 3a + 6b when a = 5 and b = –2.4. A. 0.6 B. –0.6 C. –6.4 D. 6.6

OpenStudy (anonymous):

thats it

OpenStudy (tootzrll):

for the very first equation, multiply 2 by 4 (since x equals 4). so ur equation now is 31 + 8 = 39

OpenStudy (anonymous):

if u answer all i will fan and medal

OpenStudy (tootzrll):

i can try, but fair warning it cood be wrong

OpenStudy (anonymous):

ok

OpenStudy (tootzrll):

so im working on the 0.53 + 5 one... hold on for a sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

its a got to the next 1

OpenStudy (tootzrll):

for the second one, there are 2 parts. the answer for the n= (positive) 3 is:6.59 for the second part, (the negative 3), the answer is: 3.41

OpenStudy (anonymous):

ok

OpenStudy (tootzrll):

4a+5b= is -7.5

OpenStudy (tootzrll):

9n - 10 for the fourth one

OpenStudy (tootzrll):

12m? for the fifth one

OpenStudy (tootzrll):

36 + 22d?

OpenStudy (anonymous):

number 8

OpenStudy (tootzrll):

the last one?

OpenStudy (anonymous):

yeah

OpenStudy (tootzrll):

0.6

OpenStudy (tootzrll):

do u want more help with the others?